# SPARKE: Scalable Prompt-Aware Diversity Guidance in Diffusion Models via RKE Score

### v1

@@ B.2 Proof of Proposition 2
Let $K_{Z}$ be the kernel matrix for the data $Z$ with entries $(K_{Z})_{ij}=k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})$. Let $K_{Y}$ be the kernel matrix for the conditions $Y$ with entries $(K_{Y})_{ij}=k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})$. The normalized kernel matrices are $\widetilde{K}_{Z}=K_{Z}/{\mathrm{Tr}(K_{Z})}$ and $\widetilde{K}_{Y}=K_{Y}/{\mathrm{Tr}(K_{Y})}$.
@@
$\mathrm{Tr}(K_{Z})=\sum_{i=1}^{n}k_{Z}(\bm{z}^{(i)},\bm{z}^{(i)})=\sum_{i=1}^{n}1=n$
$\mathrm{Tr}(K_{Y})=\sum_{i=1}^{n}k_{Y}(\bm{y}^{(i)},\bm{y}^{(i)})=\sum_{i=1}^{n}1=n$
@@
$\mathcal{L}_{\text{Cond-IRKE}}(\bm{z}^{(1)},\ldots,\bm{z}^{(n)};\bm{y}^{(1)},\ldots,\bm{y}^{(n)})=\|\widetilde{K}_{Z}\odot\widetilde{K}_{Y}\|_{F}^{2},$ (32)
@@
$=\frac{1}{n}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})\cdot\frac{1}{n}k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})$
$=\frac{1}{n^{2}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)}),$
@@
The squared Frobenius norm is the sum of the squares of its elements, then we have the following according to Eq. equation 32:
@@
$\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ $=\sum_{i=1}^{n}\sum_{j=1}^{n}\left(\frac{1}{n^{2}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})\right)^{2}$ (34)
$=\frac{1}{n^{4}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})^{2},$
@@
Now, we want to compute the gradient of $\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ with respect to $\bm{z}^{(n)}$ according to Eq. equation 34:
@@
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)=\frac{1}{n^{4}}\nabla_{\bm{z}^{(n)}}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})^{2}\right),$ (35)
@@
Let $S_{C}=\sum_{i=1}^{n}\sum_{j=1}^{n}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})^{2}$. We need to find $\nabla_{\bm{z}^{(n)}}S_{C}$. The terms in the sum $S_{C}$ that involve $\bm{z}^{(n)}$ (through $k_{Z}$) are those where $i=n$ or $j=n$ (or both), and the terms $k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})$ do not depend on $\bm{z}^{(n)}$. We can split the sum as follows:
$S_{C}$ $=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})^{2}\quad(\text{terms not involving }\bm{z}^{(n)})$ (36)
$+\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\quad(j=n,i\neq n)$
$+\sum_{j=1}^{n-1}k_{Z}(\bm{z}^{(n)},\bm{z}^{(j)})^{2}k_{Y}(\bm{y}^{(n)},\bm{y}^{(j)})^{2}\quad(i=n,j\neq n)$
$+k_{Z}(\bm{z}^{(n)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(n)},\bm{y}^{(n)})^{2},\quad(i=n,j=n)$
@@
Using the symmetry of the kernels, $k_{Z}(\bm{z}^{(n)},\bm{z}^{(j)})=k_{Z}(\bm{z}^{(j)},\bm{z}^{(n)})$ and $k_{Y}(\bm{y}^{(n)},\bm{y}^{(j)})=k_{Y}(\bm{y}^{(j)},\bm{y}^{(n)})$, the second and third sums are identical. Then we have:
$S_{C}$ $=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})^{2}$ (37)
$+2\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}+k_{Z}(\bm{z}^{(n)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(n)},\bm{y}^{(n)})^{2},$
Now, we compute the gradient of $S_{C}$ with respect to $\bm{z}^{(n)}$:
$\nabla_{\bm{z}^{(n)}}S_{C}=$ $\nabla_{\bm{z}^{(n)}}\left(\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(j)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(j)})^{2}\right)$ (38)
$+\nabla_{\bm{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\right)$
$+\nabla_{\bm{z}^{(n)}}\left(k_{Z}(\bm{z}^{(n)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(n)},\bm{y}^{(n)})^{2}\right),$
@@
The first term is zero because it does not depend on $\bm{z}^{(n)}$. For the second term:
$\nabla_{\bm{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\right)$ $=2\sum_{i=1}^{n-1}k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\nabla_{\bm{z}^{(n)}}\left(k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})^{2}\right)$ (39)
$=2\sum_{i=1}^{n-1}k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\left(2k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})\right)$
$=4\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\nabla_{\bm{z}^{(n)}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)}),$
@@
For the third term, since $k_{Z}(\bm{z}^{(n)},\bm{z}^{(n)})=1$ and $k_{Y}(\bm{y}^{(n)},\bm{y}^{(n)})=1$ (normalized kernels), this term is constant:
$\nabla_{\bm{z}^{(n)}}\left(k_{Z}(\bm{z}^{(n)},\bm{z}^{(n)})^{2}k_{Y}(\bm{y}^{(n)},\bm{y}^{(n)})^{2}\right)=\nabla_{\bm{z}^{(n)}}(1^{2}\cdot 1^{2})=\nabla_{\bm{z}^{(n)}}(1)=\bm{0},$ (40)
@@
$\nabla_{\bm{z}^{(n)}}S_{C}=4\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\nabla_{\bm{z}^{(n)}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)}),$ (41)
Substituting this back into the gradient of $\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ in Eq. equation 34:
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ $=\frac{1}{n^{4}}\left(4\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\nabla_{\bm{z}^{(n)}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})\right)$ (42)
$=\frac{4}{n^{4}}\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\nabla_{\bm{z}^{(n)}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)}),$
@@
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}\propto\sum_{i=1}^{n-1}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)})k_{Y}(\bm{y}^{(i)},\bm{y}^{(n)})^{2}\nabla_{\bm{z}^{(n)}}k_{Z}(\bm{z}^{(i)},\bm{z}^{(n)}).$ (43)
@@
Regarding the monotonic relationship: From Eq. equation 10, we have:
@@
We have shown that $\mathrm{Tr}(K_{Y}\odot K_{Z})=n$, since $(K_{Y}\odot K_{Z})_{ii}=k_{Y}(\bm{y}^{(i)},\bm{y}^{(i)})k_{Z}(\bm{z}^{(i)},\bm{z}^{(i)})=1\cdot 1=1$. Also, $\|\widetilde{K}_{Y}\|^{2}_{F}=\left\|K_{Y}/n\right\|^{2}_{F}=\|K_{Y}\|^{2}_{F}/n^{2}$. Substituting these into the formulation for $\mathcal{L}_{\text{Cond-RKE}}$ in Eq. equation 44:
@@
From the definition of $\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ in Eq. equation 32, we have:
@@
Let $C_{Y}=\|K_{Y}\|_{F}^{2}$. This term depends only on the conditions $Y$ and is constant with respect to $\bm{z}^{(n)}$. $C_{Y}$ is always positive since the kernel function $K_{Y}$ is positive semi-definite and not all kernel values can be 0. From the formulation for $\mathcal{L}_{\text{Cond-IRKE}}$ in Eq. equation 32, we have $\|K_{Z}\odot K_{Y}\|_{F}^{2}=n^{4}\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$. Substituting this into the formulation for $\mathcal{L}_{\text{Cond-RKE}}$:

### v2

@@ B.2 Proof of Proposition 2
Let $K_{Z}$ be the kernel matrix for the data $Z$ with entries $(K_{Z})_{ij}=k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})$. Let $K_{Y}$ be the kernel matrix for the conditions $Y$ with entries $(K_{Y})_{ij}=k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})$. The normalized kernel matrices are $\widetilde{K}_{Z}=K_{Z}/{\mathrm{Tr}(K_{Z})}$ and $\widetilde{K}_{Y}=K_{Y}/{\mathrm{Tr}(K_{Y})}$.
@@
$\mathrm{Tr}(K_{Z})=\sum_{i=1}^{n}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(i)})=\sum_{i=1}^{n}1=n$
$\mathrm{Tr}(K_{Y})=\sum_{i=1}^{n}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(i)})=\sum_{i=1}^{n}1=n$
@@
$\mathcal{L}_{\text{Cond-IRKE}}(\boldsymbol{z}^{(1)},\ldots,\boldsymbol{z}^{(n)};\boldsymbol{y}^{(1)},\ldots,\boldsymbol{y}^{(n)})=\|\widetilde{K}_{Z}\odot\widetilde{K}_{Y}\|_{F}^{2},$ (31)
@@
$=\frac{1}{n}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})\cdot\frac{1}{n}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})$
$=\frac{1}{n^{2}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)}),$
@@
The squared Frobenius norm is the sum of the squares of its elements, then we have the following according to Eq. (31):
@@
$\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ $=\sum_{i=1}^{n}\sum_{j=1}^{n}\left(\frac{1}{n^{2}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})\right)^{2}$ (33)
$=\frac{1}{n^{4}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})^{2},$
@@
Now, we want to compute the gradient of $\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ with respect to $\boldsymbol{z}^{(n)}$ according to Eq. (33):
@@
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)=\frac{1}{n^{4}}\nabla_{\boldsymbol{z}^{(n)}}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})^{2}\right),$ (34)
@@
Let $S_{C}=\sum_{i=1}^{n}\sum_{j=1}^{n}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})^{2}$. We need to find $\nabla_{\boldsymbol{z}^{(n)}}S_{C}$. The terms in the sum $S_{C}$ that involve $\boldsymbol{z}^{(n)}$ (through $k_{Z}$) are those where $i=n$ or $j=n$ (or both), and the terms $k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})$ do not depend on $\boldsymbol{z}^{(n)}$. We can split the sum as follows:
$S_{C}$ $=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})^{2}\quad(\text{terms not involving }\boldsymbol{z}^{(n)})$ (35)
$+\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\quad(j=n,i\neq n)$
$+\sum_{j=1}^{n-1}k_{Z}(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(j)})^{2}k_{Y}(\boldsymbol{y}^{(n)},\boldsymbol{y}^{(j)})^{2}\quad(i=n,j\neq n)$
$+k_{Z}(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(n)},\boldsymbol{y}^{(n)})^{2},\quad(i=n,j=n)$
@@
Using the symmetry of the kernels, $k_{Z}(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(j)})=k_{Z}(\boldsymbol{z}^{(j)},\boldsymbol{z}^{(n)})$ and $k_{Y}(\boldsymbol{y}^{(n)},\boldsymbol{y}^{(j)})=k_{Y}(\boldsymbol{y}^{(j)},\boldsymbol{y}^{(n)})$, the second and third sums are identical. Then we have:
$S_{C}$ $=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})^{2}$ (36)
$+2\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}+k_{Z}(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(n)},\boldsymbol{y}^{(n)})^{2},$
Now, we compute the gradient of $S_{C}$ with respect to $\boldsymbol{z}^{(n)}$:
$\nabla_{\boldsymbol{z}^{(n)}}S_{C}=$ $\nabla_{\boldsymbol{z}^{(n)}}\left(\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})^{2}\right)$ (37)
$+\nabla_{\boldsymbol{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\right)$
$+\nabla_{\boldsymbol{z}^{(n)}}\left(k_{Z}(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(n)},\boldsymbol{y}^{(n)})^{2}\right),$
@@
The first term is zero because it does not depend on $\boldsymbol{z}^{(n)}$. For the second term:
$\nabla_{\boldsymbol{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\right)$ $=2\sum_{i=1}^{n-1}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\nabla_{\boldsymbol{z}^{(n)}}\left(k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}\right)$ (38)
$=2\sum_{i=1}^{n-1}k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\left(2k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\right)$
$=4\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\nabla_{\boldsymbol{z}^{(n)}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}),$
@@
For the third term, since $k_{Z}(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})=1$ and $k_{Y}(\boldsymbol{y}^{(n)},\boldsymbol{y}^{(n)})=1$ (normalized kernels), this term is constant:
$\nabla_{\boldsymbol{z}^{(n)}}\left(k_{Z}(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2}k_{Y}(\boldsymbol{y}^{(n)},\boldsymbol{y}^{(n)})^{2}\right)=\nabla_{\boldsymbol{z}^{(n)}}(1^{2}\cdot 1^{2})=\nabla_{\boldsymbol{z}^{(n)}}(1)=\boldsymbol{0},$ (39)
@@
$\nabla_{\boldsymbol{z}^{(n)}}S_{C}=4\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\nabla_{\boldsymbol{z}^{(n)}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}),$ (40)
Substituting this back into the gradient of $\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ in Eq. (33):
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ $=\frac{1}{n^{4}}\left(4\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\nabla_{\boldsymbol{z}^{(n)}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\right)$ (41)
$=\frac{4}{n^{4}}\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\nabla_{\boldsymbol{z}^{(n)}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}),$
@@
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{Cond-IRKE}}\propto\sum_{i=1}^{n-1}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(n)})^{2}\nabla_{\boldsymbol{z}^{(n)}}k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}).$ (42)
@@
Regarding the monotonic relationship: From Eq. (9), we have:
@@
We have shown that $\mathrm{Tr}(K_{Y}\odot K_{Z})=n$, since $(K_{Y}\odot K_{Z})_{ii}=k_{Y}(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(i)})k_{Z}(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(i)})=1\cdot 1=1$. Also, $\|\widetilde{K}_{Y}\|^{2}_{F}=\left\|K_{Y}/n\right\|^{2}_{F}=\|K_{Y}\|^{2}_{F}/n^{2}$. Substituting these into the formulation for $\mathcal{L}_{\text{Cond-RKE}}$ in Eq. (43):
@@
From the definition of $\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$ in Eq. (31), we have:
@@
Let $C_{Y}=\|K_{Y}\|_{F}^{2}$. This term depends only on the conditions $Y$ and is constant with respect to $\boldsymbol{z}^{(n)}$. $C_{Y}$ is always positive since the kernel function $K_{Y}$ is positive semi-definite and not all kernel values can be 0. From the formulation for $\mathcal{L}_{\text{Cond-IRKE}}$ in Eq. (31), we have $\|K_{Z}\odot K_{Y}\|_{F}^{2}=n^{4}\mathcal{L}_{\text{Cond-IRKE}}(Z;Y)$. Substituting this into the formulation for $\mathcal{L}_{\text{Cond-RKE}}$:
@@
B.3 Interpretation of SPARKE
Our method, SPARKE, can be theoretically interpreted as an interacting particle system where sequential samples $\mathbf{z}^{(1)},\mathbf{z}^{(2)},...,\mathbf{z}^{(n)}$ evolve along a gradient flow to minimize an energy potential, thus improving prompt-aware diversity. This process is described by the following sequential stochastic differential equation:
$\mathrm{d}\mathbf{z}^{(n)}=\left[-\mathbf{f}(\mathbf{z}^{(n)},t^{\prime})+\frac{g^{2}(t^{\prime})}{2}\left(\nabla_{\mathbf{z}^{(n)}}\log p_{t^{\prime}}(\mathbf{z}^{(n)})-\eta\nabla_{\mathbf{z}^{(n)}}L_{\text{Cond-IRKE}}(\mathbf{z}_{1},\dots,\mathbf{z}_{n})\right)\right]\mathrm{d}t+g(t^{\prime})\mathrm{d}\mathbf{w},$ (47)
The core insight of SPARKE is to employ the Conditional Rényi Kernel Entropy (Conditional RKE) as a differentiable objective for promoting diversity. Following jalali2023information , the order-2 RKE score serves as a differentiable objective for counting the modes of a particle distribution. Furthermore, the Conditional RKE score in jalali2024conditional quantifies the internal diversity conditioning on the prompt categories. SPARKE defines a gradient flow that optimizes the Conditional RKE score, and sequentially drives samples to cover distinct modes, while the conditional mechanism ensures the prompt-aware diversity guidance potential.